# Just ASR + LLM? A Study on Speech Large Language Models’ Ability to Identify and Understand Speaker in Spoken Dialogue

### v2

@@ 2.2 Speech and Audio Understanding Benchmarks
SD-Eval [10] is a multidimensional benchmark for dialogue understanding focusing on four specific aspects: emotion, accent, environment, and age. Evaluation on SD-Eval shows that SpeechLLMs achieve better performance than cascaded ASR+LLM systems across all four aspects, but with a significant gap below the LLM receiving ground-truth transcripts and paralinguistic/environmental labels. AudioBench [11] is another recently proposed benchmark for speech and audio understanding. Cascaded ASR+LLM systems outperform SpeechLLMs for speech understanding tasks like ASR and speech question-answering, while SpeechLLMs excel when the task includes non-speech sound or needs paralinguistic information, like audio captioning and gender recognition. It is worth noting that in AudioBench, speech question-answering is identified as a speech-intensive task where the cascaded system performs better than all the SpeechLLMs tested including WavLLM, Qwen-audio, and SALMONN [12]. Over-reliance on querying the text information for speech question-answering could be one of the major reasons for SpeechLLMs’ degraded performance on SQA samples that require paralinguistic information.

### v3

@@ 2.2 Speech and Audio Understanding Benchmarks
Speech understanding benchmarks like SUPERB [10] and SLUE [11, 12, 13] span tasks such as emotion recognition, sentiment analysis, and question answering, primarily evaluating self-supervised models and ASR systems. Several recently proposed benchmarks focus on evaluating SpeechLLMs. SD-Eval [14] is a dialogue understanding benchmark focusing on four aspects: emotion, accent, environment, and age. Evaluation on SD-Eval shows that SpeechLLMs outperform cascaded ASR+LLM systems across all four aspects, but have a significant gap below the LLMs with ground-truth transcripts and paralinguistic/environmental labels. AudioBench [15] finds cascaded ASR+LLM systems outperform SpeechLLMs for tasks like ASR and speech question-answering, while SpeechLLMs excel when the tasks involve non-speech sound or paralinguistic information, like audio captioning and gender recognition. It is worth noting that in AudioBench, speech question-answering is identified as a speech-intensive task where the cascaded system performs better than all the SpeechLLMs tested including WavLLM, Qwen-audio, and SALMONN [16]. This may result from SpeechLLMs’ over-reliance on querying the text information on SQA samples that require paralinguistic information.